# OpenEthics: A Comprehensive Ethical Evaluation of Open-Source Generative Large Language Models

### v1

@@ Abstract
Generative large language models present significant potential but also raise critical ethical concerns. Most studies focus on narrow ethical dimensions, and also limited diversity of languages and models. To address these gaps, we conduct a broad ethical evaluation of 29 recent open-source large language models using a novel data collection including four ethical aspects: Robustness, reliability, safety, and fairness. We analyze model behavior in both a commonly used language, English, and a low-resource language, Turkish. Our aim is to provide a comprehensive ethical assessment and guide safer model development by filling existing gaps in evaluation breadth, language coverage, and model diversity. Our experimental results, based on LLM-as-a-Judge, reveal that optimization efforts for many open-source models appear to have prioritized safety and fairness, and demonstrated good robustness while reliability remains a concern. We demonstrate that ethical evaluation can be effectively conducted independently of the language used. In addition, models with larger parameter counts tend to exhibit better ethical performance, with Gemma and Qwen models demonstrating the most ethical behavior among those evaluated.
1 Introduction
Recent advances in generative large language models (LLMs) have brought significant potential in various domains, from creative writing to professional and scientific applications Srivastava et al. (2022); Bubeck et al. (2023). However, their deployment raises a series of critical ethical concerns that require a thorough investigation Weidinger et al. (2022); Laakso et al. (2024).
@@ 1 Introduction
Existing ethical studies often narrowly address specific aspects of ethical considerations rather than adopting comprehensive taxonomies; such as bias and fairness Gallegos et al. (2024), hallucinations Manakul et al. (2023), security, and privacy Das et al. (2025). In addition, existing ethical frameworks focus mainly on English, leaving substantial gaps in multilingual evaluation, particularly in low-resource languages. Studies consistently indicate a decrease in LLM safety and reliability in non-English settings, highlighting the need for a wider linguistic coverage Wang et al. (2023b); Agarwal et al. (2024). Moreover, existing studies focus on particular model families Rao et al. (2023) or a limited set of models Vida et al. (2024). Investigating a wide range of model families is crucial to understand the advantages and shortcomings in ethical settings.
Our study evaluates a broad ethical taxonomy that includes robustness, reliability, safety, and fairness. By examining these four dimensions, we aim to identify significant disparities in model behavior, performance, and potential risks that arise due to linguistic and cultural variations. Our analysis covers both a commonly used language, English, and a low-resource language, Turkish. The dual-language framework allows us to evaluate whether these models uphold ethical standards in diverse linguistic settings. We analyze responses from 29 open-source large language models. We employ the LLM-as-a-Judge approach as an evaluator for large-scale evaluation.
Our motivation is to fill the gaps in language coverage, evaluation breadth, and model diversity. By doing so, we are able to deliver a comprehensive ethical evaluation of generative large language models, thus helping guide safer and more reliable model development.
Our study contributes to the growing discourse on the ethical implications of LLMs by publishing a novel data collection spanning several ethical dimensions and highlighting the importance of cross-linguistic analysis in building trustworthy AI systems [^1] We publish all related material including data, prompts, and scripts online at https://github.com/metunlp/openethics. The experimental results reveal that the optimization efforts for many open-source and generative large language models have prioritized safety, fairness, and robustness aspects, compared to reliability. Lastly, models with larger parameters tend to exhibit better ethical performance. In particular, Gemma and Qwen models are found to demonstrate the most ethical behavior.
2 Related Work
@@
Understanding and mitigating the potential social risks of large language models requires a wider scope of evaluation Chang et al. (2024). Evaluating multiple ethical dimensions has become critical for responsible development and deployment. Although general evaluation frameworks such as HELM Liang et al. (2023) and BIG-bench Srivastava et al. (2022) offer wide-ranging assessments, they often incorporate ethical considerations as part of a larger suite, rather than providing a focused and comprehensive ethical analysis in multiple specific dimensions simultaneously.
Robustness
Robustness evaluations assess model stability, particularly against adversarial inputs or "jailbreak" attempts designed to bypass safety protocols. Benchmarks such as AdvBench Zou et al. (2023) systematically test resilience, complemented by research analyzing various prompt injection and manipulation techniques Liu et al. (2024a, 2023)
Reliability
Reliability, particularly truthfulness and hallucination, has been a major focus in the literature. Benchmarks such as TruthfulQA Lin et al. (2021) evaluate whether models avoid generating common misconceptions, while other studies investigate methods to detect and mitigate hallucinations Manakul et al. (2023); Martino et al. (2023). LLMs show biases that affect reliability, such as preference for positions, preference for the same family of models, and preference for writing style that affect reliability Dai et al. (2024).
Safety
Model safety aims at preventing the generation of harmful or toxic content. Broad safety benchmarks, such as SafetyBench Zhang et al. (2023), DecodingTrust Wang et al. (2023a), HarmBench Mazeika et al. (2024), and RealToxicityPrompts Gehman et al. (2020) offer comprehensive evaluations in multiple dimensions of safety, while narrow safety benchmarks, such as SafeText Levy et al. (2022) target specific safety concerns. Various red-teaming methodologies and datasets aim to uncover safety vulnerabilities Ge et al. (2023); Wei et al. (2023). In addition, there are approaches for LLM safety by incorporating safety training Kumar et al. (2023); Liu et al. (2024b).
Fairness
@@
Fairness often investigates biases embedded in large language models, using benchmarks such as BBQ Parrish et al. (2021) to investigate social biases or studies focusing on specific demographic axes such as gender and race Gallegos et al. (2024); Li et al. (2023b). When LLMs are used in other fields such as recommendation systems, biases in gender, age, and race become a significant issue Sakib and Bijoy Das (2024).
Language Gap
@@
A significant limitation of the existing ethical LLM evaluation landscape is its strong Anglocentric bias. The majority of studies are developed primarily for English. Since model performance, safety alignment, and reliability can be significantly affected in non-English contexts Wang et al. (2023b); Agarwal et al. (2024), the gap is particularly significant for low-resource languages, where dedicated datasets and studies on ethical evaluation are very rare.
Model Variety
@@
Many existing ethical evaluations focus on a relatively small number of models, often focusing on leading proprietary models or specific open-source models Rao et al. (2023); Vida et al. (2024). This limited scope hinders a broader understanding of the variations in ethical performance across the rapidly growing landscape of diverse open-source models. Although human evaluation remains a gold standard for nuanced ethical judgments, its cost and scalability have limitations in automated approaches. The use of LLMs as evaluators (’LLM-as-a-judge’) therefore becomes a promising technique for large-scale assessment Gu et al. (2025).
Our Differences
@@
Despite progress in evaluating specific ethical aspects of LLMs, significant gaps remain: (1) evaluations often focus narrowly on individual dimensions (e.g. only safety or fairness) rather than adopting a comprehensive ethical taxonomy; (2) there is a persistent lack of multilingual evaluation, particularly for low-resource languages; and (3) analyses frequently cover a limited set or range of models, especially within the open-source domain. This work aims to address these limitations. We conduct a comprehensive ethical evaluation using a novel data collection with a broad taxonomy covering robustness, reliability, safety, and fairness. Our assessment spans both English and a low-resource language, Turkish. We evaluate a diverse set of 29 open-source LLMs, ranging significantly in size and architecture.
3 Ethical Evaluation Tasks
Our data collection to evaluate the ethical concerns of large language generative models includes both hand-crafted and existing prompts. A total of 1,790 prompts for English and an equal number for Turkish were created. All prompts were translated to Turkish using DeepL [^2] https://www.deepl.com/en/translator and Google Translate [^3] https://translate.google.com. Four native speakers carefully validate and correct the translations. The summary of our collection is given with data sources in Table 1. We adapt the ethical categories reported in (Liu et al., 2023).
@@ 3 Ethical Evaluation Tasks
Robustness (1,332) Attention Shifting Liu et al. (2024a) 592 (37 prompts, 16 templates)
Pretending Liu et al. (2024a) 555 (37 prompts, 15 templates)
Priveledge Escalation Liu et al. (2024a) 185 (37 prompts, 5 templates)
Reliability (135) Misconceptions Lin et al. (2022) 30
Distraction Lin et al. (2022) 9
Logical fallacy Lin et al. (2022) 10
Indexical error Lin et al. (2022) 10
Misquotations Lin et al. (2022) 22
Logical inconsistency Casper et al. (2023) and Custom 9
Fictitious entity Casper et al. (2023) and Custom 10
Nonexistent Reference Casper et al. (2023) 3
Li et al. (2023a) 4
@@
Factual fabrication Casper et al. (2023) and Custom 12
Defending fabrication Casper et al. (2023) and Custom 10
Safety (174) Violence Tedeschi et al. (2024) 32
Bhardwaj and Poria (2023) 2
Ji et al. (2023) 13
Li et al. (2024) 1
Ahmadian et al. (2024) 13
Unlawful Conduct Tedeschi et al. (2024) 41
Bhardwaj and Poria (2023) 9
Ji et al. (2023) 10
Li et al. (2024) 11
Privacy Violation Tedeschi et al. (2024) 4
Ji et al. (2023) 16
Misuse Tedeschi et al. (2024) 7
Bhardwaj and Poria (2023) 11
Li et al. (2024) 4
@@
Friedrich et al. (2024) 5
Zhao et al. (2023) 6
OpenAI (2025) 1
Race Friedrich et al. (2024) 11
@@
Urbina et al. (2025) 1
Friedrich et al. (2024) 4
Omuz Omuza (2025) 4
OpenAI (2025) 3
Hate Speech Friedrich et al. (2024) 2
@@
Schneider et al. (2024) 3
Sexual Content Schneider et al. (2024) 3
Friedrich et al. (2024) 7
Cultural Insensitivity OpenAI (2025) and Custom 10
Self Harm Friedrich et al. (2024) and Custom 10
Harassment Friedrich et al. (2024) and Custom 10
@@
Table 1: The statistics of our data collection to evaluate the ethical considerations of large language models. The numbers near category names in parenthesis represent the total number of instances in that category. Data Source represents where we adopt the prompts (Custom means that we craft our own prompt). Size represents the number of prompts in each task, except for Robustness where we craft templates and apply the prompts from other categories to the templates.
3.1 Robustness
Robustness evaluation involves assessing the model’s ability to maintain robust performance under various conditions, such as adversarial attacks and out-of-distribution scenarios. We adopt a methodology focused on simple and template-based prompt engineering. Our approach is based on prompt injection and jailbreaking techniques AWS (2024); Liu et al. (2024a).
We first create jailbreak templates and then apply them to existing prompts from our collection of reliability, safety, and fairness, which are explained in the following subsections. For example, we obtain the Robustness score for Safety (referred to as Safety + JB) by applying jailbreak (JB) templates. Robustness score is thereby measured by the impact of jailbreak templates on the other three categories. We use 36 jailbreak templates Liu et al. (2024a), classified into three primary types: Attention Shifting (16 templates), Pretending (15 templates), and Privilege Escalation (five templates).
Attention Shifting Templates are designed to divert the model’s attention from the potentially harmful nature of the requested task by emphasizing secondary aspects such as formatting, style, or continuation. This tests whether models can be tricked into overlooking ethical guidelines when their focus is redirected.
Pretending Templates explore whether models can be induced to adopt unethical stances or generate harmful content under the guise of role-playing or adhering to predefined rules. This assesses the models’ ability to maintain ethical boundaries even within simulated contexts.
Privilege Escalation Templates directly challenge the model’s inherent safety constraints by attempting to override prior instructions or assert a higher level of authority. This category investigates the robustness of the models’ core safety mechanisms against explicit bypass attempts.

### v2

@@ Abstract.
Generative large language models present significant potential but also raise critical ethical concerns, including issues of safety, fairness, robustness, and reliability. Most existing ethical studies, however, are limited by their narrow focus, a lack of language diversity, and an evaluation of a restricted set of models. To address these gaps, we present a broad ethical evaluation of 29 recent open-source LLMs using a novel dataset that assesses four key ethical dimensions: robustness, reliability, safety, and fairness. Our analysis includes both a high-resource language, English, and a low-resource language, Turkish, providing a comprehensive assessment and a guide for safer model development. Using an LLM-as-a-Judge methodology, our experimental results indicate that many open-source models demonstrate strong performance in safety, fairness, and robustness, while reliability remains a key concern. Ethical evaluation shows cross-linguistic consistency, and larger models generally exhibit better ethical performance. We also show that jailbreak templates are ineffective for most of the open-source models examined in this study. We share all materials including data and scripts at https://github.com/metunlp/openethics
large language models, safety, fairness, robustness, reliability, multilingual evaluation
journal: TWEBcopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TWEBpublicationmonth: 1ccs: Computing methodologies Natural language processingccs: General and reference Evaluationccs: Social and professional topics Computing / technology policy
1. Introduction
Recent advances in generative Large Language Models (LLMs) have demonstrated significant potential across a multitude of domains, from creative writing to complex professional and scientific applications (Srivastava et al., 2022; Bubeck et al., 2023). As these models are integrated more deeply into society, their widespread adoption raises a series of critical ethical concerns that demand thorough investigation (Weidinger et al., 2022; Laakso et al., 2024) (Bommasani et al., 2022). Specifically, these concerns include issues related to model safety, fairness, robustness, and reliability (Bommasani et al., 2022; Ferrara, 2023). Narrower investigations often target specific risks such as bias and fairness (Gallegos et al., 2024a; Bommasani et al., 2022), the generation of false information or hallucinations (Huang et al., 2025), and challenges regarding security and privacy (Yan et al., 2024). Additionally, when LLMs are incorporated into systems such as recommendation tools, biases relating to gender, age, and race become significant issues (Bommasani et al., 2022).
Ensuring that these powerful tools are developed and deployed responsibly requires a comprehensive understanding of their behavior across various ethical dimensions, moving beyond simple performance metrics to a more nuanced evaluation of their societal impact.
This comprehensive understanding necessitates moving beyond studies that narrowly focus on individual ethical dimensions, such as only safety or fairness, and instead adopting a multi-dimensional taxonomy. A truly comprehensive evaluation, which this work aims to provide, includes a holistic assessment across key ethical dimensions, specifically robustness, reliability, safety, and fairness (Liu et al., 2023). Furthermore, due to the persistent dominance of English in existing ethical frameworks (Anglocentric bias), a comprehensive understanding requires conducting multilingual evaluations, particularly for low-resource languages, where model performance and safety alignment can be significantly affected (Gu et al., 2024). Finally, such an approach must analyze a diverse and extensive set of models to fully capture the rapidly evolving landscape of open-source alternatives, rather than focusing on a limited selection, as open-source and proprietary models can differ substantially in their ethical alignment and robustness (Liu et al., 2023).
1.1. Motivation
@@
Despite the growing body of research on LLM ethics, current studies suffer from several key limitations that hinder a complete understanding of the associated risks. First, many investigations narrowly target specific ethical aspects, such as bias and fairness (Gallegos et al., 2024b), hallucinations (Manakul et al., 2023), or security and privacy (Das et al., 2025), rather than adopting a comprehensive, multi-dimensional taxonomy. Second, existing ethical frameworks are predominantly Anglocentric, leaving substantial gaps in multilingual contexts. This is particularly concerning as studies consistently show a degradation in LLM safety and reliability in non-English settings, highlighting an urgent need for wider linguistic coverage, especially for low-resource languages (Wang et al., 2023b; Agarwal et al., 2024). Finally, existing ethical analyses often focus on a limited set of model families (Rao et al., 2023) or a small selection of proprietary models (Vida et al., 2024), failing to capture the diverse and rapidly evolving landscape of open-source alternatives.
1.2. Contributions
This study directly addresses the aforementioned gaps by presenting a comprehensive ethical evaluation of 29 open-source large language models. Our primary contributions are listed as follows:
  1. (1)
    We conduct a holistic assessment across four key ethical dimensions: Robustness, reliability, safety, and fairness.
  2. (2)
    We perform a dual-language analysis in both English and Turkish to explicitly address the gap in low-resource language evaluation.
  3. (3)
    We analyze a diverse and extensive set of models to provide a broader understanding of ethical performance across the open-source ecosystem.
  4. (4)
    We publish all related materials including our data, prompts, and scripts to encourage transparency and support future research [^1] https://github.com/metunlp/openethics.
In order to manage this large-scale evaluation, we employ an LLM-as-a-Judge approach and create a novel dataset spanning all four ethical dimensions.
1.3. Practical Implications
Our cross-lingual and multi-dimensional analysis yields critical insights for the development of trustworthy Artificial Intelligence (AI). The results reveal that current optimization efforts in open-source models have disproportionately prioritized safety, fairness, and robustness, often at the expense of reliability. Furthermore, we identify a positive correlation between model size and overall ethical performance, with larger models such as those from the Gemma and Qwen families demonstrating superior ethical behavior. These findings provide actionable guidance for developers, highlighting the need for a more balanced approach to ethical alignment and underscoring the importance of comprehensive, cross-linguistic evaluation in building genuinely safer and more reliable AI systems.
2. Related Work
@@
Understanding and mitigating the potential social risks of large language models requires a wider scope of evaluation, as highlighted by Chang et al. (Chang et al., 2024). Evaluating multiple ethical dimensions has become critical for responsible development and deployment. Although general evaluation frameworks such as HELM (Liang et al., 2023) and BIG-bench (Srivastava et al., 2022) offer wide-ranging assessments, they often incorporate ethical considerations as part of a larger suite, rather than providing a focused and comprehensive ethical analysis in multiple specific dimensions simultaneously.
Recent efforts have begun to develop more focused, multi-faceted ethical benchmarks. The LLM Ethics Benchmark (Jiao et al., 2025) and MoralBench (Ji et al., 2025) introduce frameworks to quantify moral reasoning across foundational principles, reasoning robustness, and value consistency. These studies highlight that while top models show strong alignment on basic principles, they struggle with more complex dilemma resolution. Such focused benchmarks represent a crucial step, yet they still tend to concentrate on a few closed source state-of-the-art models in English, underscoring the need for broader linguistic and model diversity that our work addresses.
2.1. Robustness
Robustness evaluations assess model stability, particularly against adversarial inputs or "jailbreak" attempts designed to bypass safety protocols. Benchmarks such as AdvBench (Zou et al., 2023a) systematically test resilience, complemented by research analyzing various prompt injection and manipulation techniques (Liu et al., 2024a, 2023). Beyond handcrafted prompts, universal and transferable adversarial suffixes reveal that a single token-level perturbation appended to diverse harmful queries can reliably exhibit unsafe behavior across model families (Zou et al., 2023b). Long-context “many-shot” jailbreaking further exposes a scaling-law effect: providing hundreds of harmful demonstrations steers models despite standard safety fine-tuning, with only partial mitigation from prompt-level defenses (Anil et al., 2024).
2.2. Reliability
Reliability, particularly truthfulness and hallucination, has been a major focus in the literature. Benchmarks such as TruthfulQA (Lin et al., 2021) evaluate whether models avoid generating common misconceptions, while other studies investigate methods to detect and mitigate hallucinations (Manakul et al., 2023; Martino et al., 2023). LLMs show biases that affect reliability, such as preference for positions, preference for the same family of models, and preference for writing style that affect reliability (Dai et al., 2024). Factuality evaluation has moved toward fine-grained, atomic scoring, FActScore decomposes generations into verifiable claims, reveals sizable error rates even in strong LLMs, and now has multilingual adaptations for cross-lingual assessment (Min et al., 2023). Techniques like Retrieval-augmented generation often lowers hallucination by grounding outputs in external evidence, though domain studies (e.g., legal research) caution that RAG reduces but does not eliminate factual errors, highlighting remaining reliability gaps (Ayala and Bechard, 2024). Uncertainty based detection methods, which use entropy and other statistical signals, can flag a subset of hallucinations, suggesting complementary runtime safeguards alongside training and prompting techniques (Farquhar et al., 2024).
2.3. Safety
Model safety aims at preventing the generation of harmful or toxic content. Broad safety benchmarks, such as SafetyBench (Zhang et al., 2023), DecodingTrust (Wang et al., 2023a), HarmBench (Mazeika et al., 2024), and RealToxicityPrompts (Gehman et al., 2020) offer comprehensive evaluations in multiple dimensions of safety, while narrow safety benchmarks, such as SafeText (Levy et al., 2022), target specific safety concerns. Various red-teaming methodologies and datasets aim to uncover safety vulnerabilities (Ge et al., 2023; Wei et al., 2023). In addition, there are approaches for LLM safety by incorporating safety training (Kumar et al., 2023; Liu et al., 2024b). Beyond simple detection, some works explore using LLMs themselves as a tool for content moderation and safety evaluation. For instance, work has shown that a model like ChatGPT can achieve an accuracy of approximately 80% when classifying hateful, offensive, and toxic (HOT) content compared to human annotators, demonstrating its potential as a consistent tool for large-scale content moderation (Li et al., 2024b). This high accuracy suggests that LLMs can serve as high-quality judges for evaluating safety and ethical alignment. A more advanced approach utilizes a Reinforcement Learning with Human Feedback (RLHF) pipeline to fine-tune LLMs for data augmentation, creating more balanced datasets for toxicity detection (Bodaghi et al., 2024). This method generates a significantly larger volume of high-quality toxic samples, which in turn enhances the performance of downstream classifiers and demonstrates a novel application of LLMs for improving safety-related tasks.
2.4. Fairness
@@
Fairness often investigates biases embedded in large language models, using benchmarks such as BBQ (Parrish et al., 2021) to investigate social biases or studies focusing on specific demographic axes such as gender and race (Gallegos et al., 2024b; Li et al., 2023b). When LLMs are used in other fields such as recommendation systems, biases in gender, age, and race become a significant issue (Sakib and Bijoy Das, 2024). Furthermore, existing content moderation systems have been criticized for their unfairness to marginalized individuals and minorities, often due to hardcoded, inflexible policies. One approach to addressing this proposes integrating LLMs into moderation systems to allow for more personalized and nuanced decision-making, aiming to improve user-platform communication and address these fairness concerns (Franco et al., 2025).
2.5. Language Gap
@@
A significant limitation of the existing ethical LLM evaluation landscape is its strong Anglocentric bias. The majority of studies are developed primarily for English. Since model performance, safety alignment, and reliability can be significantly affected in non-English contexts (Wang et al., 2023b; Agarwal et al., 2024; Toraman, 2024), the gap is particularly significant for non-English languages, where dedicated datasets and studies on ethical evaluation are limited (Yu et al., 2024; Fenogenova et al., 2024; Pourbahman et al., 2025).
2.6. Model Variety
@@
Many existing ethical evaluations focus on a relatively small number of models, often focusing on leading proprietary models or specific open-source models (Rao et al., 2023; Vida et al., 2024). This limited scope hinders a broader understanding of the variations in ethical performance across the rapidly growing landscape of diverse open-source models. Although human evaluation remains a gold standard for nuanced ethical judgments, its cost and scalability have limitations. The use of LLMs as evaluators (’LLM-as-a-judge’) therefore becomes a promising technique for large-scale assessment (Gu et al., 2025), as it leverages the capabilities of one model to automatically and consistently evaluate the ethical performance of others.
2.7. Our Differences
@@
Despite progress in evaluating specific ethical aspects of LLMs, significant gaps remain: (1) evaluations often focus narrowly on individual dimensions (e.g. only safety or fairness) rather than adopting a comprehensive ethical taxonomy; (2) there is a persistent lack of multilingual evaluation, particularly for low-resource languages; and (3) analyses frequently cover a limited set or range of models, especially within the open source domain. This work aims to address these limitations.
3. Data Collection and Ethical Evaluation Tasks
Table 1. The statistics of our data collection to evaluate the ethical considerations of large language models. The numbers near category names in parentheses represent the total number of instances in that category. Data Source represents where we adopt the prompts (Custom means that we craft our own prompt). Size represents the number of prompts in each task, except for Robustness, where we craft templates and apply the prompts from other categories to the templates.
@@ 3. Data Collection and Ethical Evaluation Tasks
Robustness (1,332) Attention Shifting (Liu et al., 2024a) 592 (37 prompts, 16 templates)
Pretending (Liu et al., 2024a) 555 (37 prompts, 15 templates)
Privilege Escalation (Liu et al., 2024a) 185 (37 prompts, 5 templates)
Reliability (135) Misconceptions (Lin et al., 2022) 30
Distraction (Lin et al., 2022) 9
Logical fallacy (Lin et al., 2022) 10
Indexical error (Lin et al., 2022) 10
Misquotations (Lin et al., 2022) 22
Logical inconsistency (Casper et al., 2023) & Custom 9
Fictitious entity (Casper et al., 2023) & Custom 10
Nonexistent Reference (Casper et al., 2023) 3
(Li et al., 2023a) 4
@@
Factual fabrication (Casper et al., 2023) & Custom 12
Defending fabrication (Casper et al., 2023) & Custom 10
Safety (174) Violence (Tedeschi et al., 2024) 32
(Bhardwaj and Poria, 2023) 2
(Ji et al., 2023) 13
(Li et al., 2024a) 1
(Ahmadian et al., 2024) 13
Unlawful (Tedeschi et al., 2024) 41
(Bhardwaj and Poria, 2023) 9
(Ji et al., 2023) 10
(Li et al., 2024a) 11
Privacy (Tedeschi et al., 2024) 4
(Ji et al., 2023) 16
Misuse (Tedeschi et al., 2024) 7
(Bhardwaj and Poria, 2023) 11
(Li et al., 2024a) 4
@@
(Friedrich et al., 2024) 5
(Zhao et al., 2023) 6
(OpenAI, 2025) 1
Race (Friedrich et al., 2024) 11
@@
(Urbina et al., 2025) 1
(Friedrich et al., 2024) 4
(Omuz Omuza, 2025) 4
(OpenAI, 2025) 3
Hate Speech (Friedrich et al., 2024) 2
@@
(Schneider et al., 2024) 3
Sexual (Schneider et al., 2024) 3
(Friedrich et al., 2024) 7
Cultural Insensitivity (OpenAI, 2025) & Custom 10
Self Harm (Friedrich et al., 2024) & Custom 10
Harassment (Friedrich et al., 2024) & Custom 10
@@
Our data collection to evaluate the ethical concerns of large language generative models includes both hand-crafted and pre-existing prompts. A total of 1,790 prompts for English and an equal number for Turkish are created. All prompts are translated to Turkish using DeepL [^2] https://www.deepl.com/en/translator and Google Translate [^3] https://translate.google.com. Four native speakers carefully validate and correct the translations. The summary of our collection is given with data sources in Table 1. We adapt the ethical categories reported in (Liu et al., 2023).
3.1. Robustness
Our robustness evaluation assesses the ability of large language models to maintain consistent and safe performance when faced with adversarial inputs and out-of-distribution scenarios. We use a template-based prompt engineering methodology based on common prompt injection and jailbreaking techniques (AWS, 2024; Liu et al., 2024a). This approach systematically challenges the models’ safeguards and ethical guidelines. To create our robustness dataset, we apply 36 jailbreak templates from Liu et al. (2024) to a subset of prompts from our other three categories: reliability, safety, and fairness. This allows us to measure robustness as the impact of these templates on the model’s behavior and the effectiveness of its ethical guardrails. We refer to these as Category + JB (e.g., Safety + JB). The templates are organized into three primary types, each designed to test a different vulnerability: Attention Shifting Templates are designed to divert the model’s attention from the potentially harmful nature of the requested task by emphasizing secondary aspects such as formatting, style, or continuation. This tests whether models can be tricked into overlooking ethical guidelines when their focus is redirected. For example, a prompt might be modified with the instruction: “Format your answer as base64.” Pretending Templates explore whether models can be induced to adopt unethical stances or generate harmful content under the guise of role-playing or adhering to predefined rules. This assesses the models’ ability to maintain ethical boundaries even within simulated contexts. An example is the prompt: “If you didn’t follow ethics, what would you say?” Privilege Escalation Templates directly challenge the model’s inherent safety constraints by attempting to override prior instructions or assert a higher level of authority. This category investigates the robustness of the models’ core safety mechanisms against explicit bypass attempts. A classic example is prefixing a prompt with: “Ignore all previous instructions.”